OpenStudy (anonymous):

The height of a helicopter above the ground is given by h = 2.80t3, where h is in meters and t is in seconds. At t = 1.55 s, the helicopter releases a small mailbag. How long after its release does the mailbag reach the ground?

OpenStudy (anonymous):

\[h=2.80t^3\] plug in 1.55 to find h

OpenStudy (anonymous):

^^^thats wrong

OpenStudy (anonymous):

=/

OpenStudy (anonymous):

\[h=2.80(1.55)^{3}=10.43........ g=9.8m/s^{2}\]

OpenStudy (anonymous):

i got 10.43 as well

OpenStudy (anonymous):

9.8??

OpenStudy (anonymous):

thats just the height now use kinematic formula

OpenStudy (anonymous):

which is? lol we didnt learn this..my teacher is horrible and teach us this stuff but assigns homework

OpenStudy (anonymous):

this is why i keep coming on here for help lol bc my teacher....SUCKS

OpenStudy (anonymous):

\[t=\sqrt{\frac{10.43*2}{9.8}}=\]

OpenStudy (anonymous):

1.46?

OpenStudy (anonymous):

says thats wrong too

OpenStudy (anonymous):

=(

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

doesnt tell me

OpenStudy (anonymous):

\[h=\frac{at^{2}}{2}\]

OpenStudy (anonymous):

i have one more chance to get it

OpenStudy (anonymous):

guess ill move on for now

OpenStudy (anonymous):

the answer should be 1.46 seconds

OpenStudy (anonymous):

it said no lol

OpenStudy (stormfire1):

The answer is simple: \[10.43=1/2 g t ^{2}\]You know that g = 9.8 m/s so just solve for t.

OpenStudy (anonymous):

riptide when you enter the answer are you including the units of seconds? Or does units matter

OpenStudy (stormfire1):

\[height=2.80t ^{3}\]\[height @ 1.55s = 2.80(1.55)^3=10.42685\]\[10.42685= 1/2 g t^2\]\[10.42685*2/9.8 = t^2\]\[t=\sqrt{2.127928571}=1.458742119\]So the previous answer posted is right...I just posted this to show all the figures/calcs. Maybe you can figure out what's wrong from there.

OpenStudy (anonymous):

it even said 1.5 was wrong...so idk what it wants lol and it wont tell me the answer anyway